A large tank is filled to capacity with 500 gallons of pure water. Brine containing 2 pounds of salt per gallon is pumped into the tank at a rate of 5 gal/min. The well-mixed solution is pumped out at the same rate. Find the number A(t) of pounds of salt in the tank at time t.

A(t) = ____ lb
1
Expert's answer
2020-06-22T18:13:22-0400
"{dA\over dt}=R_{in}-R_{out}"

"R_{in}=(2lb/gal)\cdot(5gal/min)=10lb/min"

"R_{out}=({A(t)\over 500}lb/gal)\cdot(5gal/min)={A(t)\over 100}lb/min"

Hence


"{dA\over dt}=10-{A(t)\over 100}"

"{dA\over 10-{A\over 100}}=dt""\ln\big|10-{A\over 100}\big|=-0.01t+\ln c"

"10-{A\over 100}=ce^{-{t\over 100}}"


"A=1000-100ce^{-{t\over 100}}"

Apply the initial conditions


"A(0)=0=1000-100ce^{-{0\over 100}}=>c=10"

"A(t)=1000-1000e^{-{t\over 100}}\ lb"